OpenStudy (anonymous):

find the 25th term of the following sequence 22,26,30...

OpenStudy (anonymous):

this is an arithmetic sequence: An=A1+(n-1)d your An= A25, A1(first term)=22 your n=25 and your d=4 try to solve it now :)

OpenStudy (anonymous):

so would it look like 22=25+(n-1)4

OpenStudy (anonymous):

no you mixed them up.. A25=22+(25-1)4---> A25=22+(24x4)

OpenStudy (anonymous):

and than i just solve it like that?

OpenStudy (anonymous):

yeah :) tell me what you got though

OpenStudy (anonymous):

right now i have a25=118 what do i do now

OpenStudy (anonymous):

thats the answer, a25 is the 25th term :)

OpenStudy (anonymous):

ohh thank you so much

OpenStudy (anonymous):

nice team work :)